You are given a stock of female Drosophila that appears wild type. You cross them with wild-type males and collect progeny, recording the following results: 1,000 wild-type females, 450 wild-type males, and 50 white-eyed males. (Note the male to female ratio.)

How many genes are assorting? What phenotypes do they control? Is linkage indicated? If so, what is the map distance?
Answer #1

Representation:

Considering trait of eye color:

Wildtype: W (dominant)

  • Male= XWY
  • Female= XW XW (Homozygous), XW Xw (Heterozygous)

White eye: w (recessive)

  • Male= XwY
  • Female= Xw Xw

Cross: Wildtype (heterozygous) Female with wildtype male:

  • XW Xw x XWY

F1

Female

Male

XW

Xw

F2

XW

XW XW

XW Xw

Female

Y

XWY

XwY

Male

  • XW XW = Wildtype female (Homozygous) = 25% =1/4
  • XW Xw = Wildtype female (Heterozygous)= 25%=1/4
  • XWY = Wildtype male = 25%=1/4
  • XwY = White eyed males = 25%=1/4

Phenotype ratio of wild type female: Wildtype male: White eyed males

= 2:1:1

Wildtype: mutant = 3:1

Thus,

  • There is one gene with two alleles (dominant for wildtype eye color and recessive for white color eye) which are assorting.
  • There independent assortment and no linkage.